# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v1

@@ 3.1 Problem Formulation
We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points. For simplicity, we use $[m]$ to denote the set $\{1,2,\ldots,m\}$. For any $i\in[m+1]$, let $q_{i}$ be the data distribution at time $t_{i}$ over $\mathbb{R}^{d}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in[m+1]}$, where $\mathbf{X}_{t_{i}}=[\bm{x}_{i}(k)]_{k\in[n_{i}]}$ is the expression matrix at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$. This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
@@
A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, define
$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),\quad u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (7)
@@
where the conditioning variable is selected as $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 6. It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$. To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
@@
Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and integrate the biological prior knowledge that can be inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension. However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 3(a) in Appendix G.1 for an illustration).
3.2 Introducing Spatial Priors & Transitional Plausibility
@@
To faithfully model the spatial context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the transitional plausibility between locations and cell states at different time points.
@@
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affects their states in a similar manner and results in local smoothness of cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals. Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
@@
$\mathrm{SS}(c_{i},c_{j})=\bigg\|\>\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\>\bigg\|_{2}^{2},$ (8)
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ is the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile of cell $c$.
@@
Specifically, we can represent each cell $c_{i}$ by a vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 7 for an illustration). We define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2},$ (9)
3.3 ContextFlow: CFM with Context-Aware OT Couplings
@@
Transitional Plausibility Matrix. First, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, let $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ be the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}\left(c_{i}(k),c_{i+1}(l)\right)+(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right),$ (10)
@@
where $\lambda\in[0,1]$ is a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
@@
Prior-Regularized OT Couplings. The transitional plausibility matrices capture our spatially informed prior on cell-cell transitions between consecutive time points, which can naturally be incorporated in the EOT formulation (Equation 6) to promote couplings that maintain the structural and functional properties of the tissue organization. We propose two techniques for prior integration:
@@
Prior-Aware Cost Matrix (PACM). Consider the empirical counterpart of Equation 6 with respect to time $t_{i}$ and time $t_{i+1}$. Our first approach incorporates the transitional plausibility matrix directly into the transport cost:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\underbrace{\Big[\alpha\cdot\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)\cdot\big[\mathbf{M}_{i,i+1}\big]_{kl}\Big]}_{\text{Prior-Aware Cost Function}}-\epsilon\sum_{k,l}\Pi_{kl}\log\Pi_{kl},$ (11)
@@
where the transport plan $\Pi$ satisfies the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 11 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This aligns with our assumption that such transitions are implausible.
@@
Prior-Aware Entropy Regularization (PAER). While the prior-aware cost matrix approach penalizes couplings in accordance with our spatial priors, it defines a different OT problem characterized by a modified cost function. Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (Proposition 1 and Corollary 1 in the Appendix C). Besides, selecting an appropriate $\alpha$ in Equation 11 introduces an additional layer of tuning, increasing computational overhead. Therefore, we propose a second approach to integrate the biological priors without introducing additional hyperparameters:
@@
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}$ $\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\log(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl})}_{\text{Prior-Aware Entropy Regularization}},$ (12)
@@
where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\exp(-[\mathbf{M}_{i,i+1}]_{kl})/\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})$ denotes the prior joint probability matrix induced by $\mathbf{M}_{i,i+1}$. Intuitively, the lower the cost $\mathbf{M}{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}{i,i+1}]_{kl}$, reflecting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. The entropy regularization term in Equation 12 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
ContextFlow. Finally, we apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problem in Equation 11 or Equation 12 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 7. The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings (ContextFlow), is detailed in Algorithm 1 in Appendix D.
In particular, to apply the Sinkhorn algorithm to solve our prior-aware entropy regularization problem in Equation 12, we make use of the following theorem, a generalized result of Peyré et al. (2019).
Theorem 1.
$\mathbf{\Pi}^{*}_{\mathrm{CTF-H}}={\arg\!\min}_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}$ $\sum_{k,l}\Pi_{kl}C_{kl}+\epsilon\sum_{k,l}\Pi_{kl}\log(\Pi_{kl}/{M}_{kl}),$
Theorem 1, proven in Appendix B, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 16. When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with Entropic-OT (Section 2.3). This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
4 Experiments
Datasets. We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021). For all the datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors. The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles. Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
Baselines & Metrics. We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several baselines. As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings. We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors. For evaluation, we employed 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, and metrics such as MMD and Energy Distance for statistical fidelity. Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Exact metric definitions are present in the Appendix F. All reported metrics are averaged across $10$ runs.
Sampling. A trained velocity field can be evaluated through the samples it generates. We consider two variants. Initial value problem sampling (IVP) integrates the learned gradient starting from the first observed batch of cells and evolves them toward a later time point. IVP provides the most comprehensive evaluation of flow quality, as errors can accumulate across steps. In contrast, next-step sampling (Next Step) integrates the gradient only from the most recently observed batch of cells, thus limiting error propagation but providing a less stringent test of long-term trajectory fidelity.
@@ 4 Experiments
4.1 Axolotl Brain Regeneration
We first evaluate ContextFlow on longitudinal Stereo-seq spatial transcriptomic data coming from a post-traumatic brain regeneration study of the Salamander (axolotl telencephalon) species (Wei et al., 2022). The dataset contains samples from five developmental stages, with replicates collected from different individual organisms at each stage. For our CTF-C method, we present the best ablated $\alpha$ in the main text, with full ablation results across different $\alpha$ values provided in Appendix H.
For extrapolation, we evaluate generation on the last holdout time point, representing the most challenging test of generalizability for the velocity fields $u_{\theta}$, as it lies outside the training time horizon. As shown in Table 2, CTF-H again consistently achieves the best overall performance, particularly under IVP-Sampling, where errors are most likely to accumulate. This result further reinforces the robustness and reliability of CTF-H across the entire sampling horizon. Finally, Figure 3 (Appendix G.1) demonstrates that incorporating spatial priors enables ContextFlow to produce substantially fewer biologically implausible couplings compared to its context-free counterpart.

### v2

@@ 3.1 Problem Formulation
We focus on the task of inferring spatiotemporal trajectories, i.e., inferring the dynamic evolution of the cell states across time from spatially resolved gene expression data. Let $0=t_{1}<t_{2}<\ldots<t_{m+1}=1$ be a sequence of normalized time points.For any $i\in\{1,2,\ldots,m+1\}$, let $q_{i}$ be the data distribution over $\mathbb{R}^{d}$ at time point $t_{i}$. Given $\{\mathbf{X}_{t_{i}}\}_{i\in\{1,2,\ldots,m+1\}}$, where $\mathbf{X}_{t_{i}}=\{\bm{x}_{i}(k)\}_{k=1}^{n_{i}}$ is the gene expressions at time $t_{i}$ consisting of $n_{i}$ snapshot data sampled from $q_{i}$, the objective is to learn a neural velocity vector field $u_{\theta}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ to faithfully characterize the temporal evolution of spatially resolved tissues over time, such that the induced probability path $p_{t}$ can describe the state of each cell at time $t\in[0,1]$. This task can be viewed as a continuous temporal generalization of the pairwise generative modeling task described in Section 2.1.
@@
A promising candidate solution is conditional flow matching with entropic OT couplings (EOT-CFM), by targeting linear conditional velocity vector fields for each pair of consecutive time points. Specifically, for any $t\in[0,1]$ satisfying $t\in[t_{i},t_{i+1}]$, the conditional velocity field $u_{t}(\bm{x}|\bm{z})$ and its associated probability path $p(\bm{x}|\bm{z})$ are defined as:
$p(\bm{x}|\bm{z})=\mathcal{N}\left(\frac{t_{i+1}-t}{t_{i+1}-t_{i}}\bm{x}_{i}+\frac{t-t_{i}}{t_{i+1}-t_{i}}\bm{x}_{i+1},\sigma^{2}\mathbf{I}\right),$
$u_{t}(\bm{x}|\bm{z})=\frac{\bm{x}_{i+1}-\bm{x}_{i}}{t_{i+1}-t_{i}},$ (8)
@@
where the conditioning variable $\bm{z}=(\bm{x}_{i},\bm{x}_{i+1})$, and $p(\bm{z})$ is the joint probability measure with marginals $q_{i}$ and $q_{i+1}$ corresponding to the EOT coupling $\pi^{*}_{\mathrm{eot}}(\epsilon)$ defined in Equation 7. It can be easily verified that the above construction satisfies the boundary condition at each time point $t_{i}$. To train $u_{\theta}$, we can randomly sample a mini-batch of data at each time, run the Sinkhorn algorithm (Cuturi, 2013) to obtain the entropic OT couplings for each consecutive pair, and iteratively update the model weights $\theta$ using stochastic gradient descent with CFM regression loss (Equation 3).
@@
Despite their enhanced ability to model system dynamics, state-of-the-art OT-CFM frameworks lack provisions to fully exploit the contextual richness and to integrate biological prior knowledge inferred from other associated data modalities. Existing approaches can generate statistically optimal trajectories by targeting probability paths induced by (entropic) OT couplings along the temporal dimension. However, they may overlook important functional or structural prior information, leading to biologically implausible trajectories (see Figure 4(a) in Appendix G.1 for an illustration).
3.2 Spatial Priors and Transitional Plausibility
@@
To faithfully model the context and cellular organization of spatial omics data, we introduce two types of spatial priors and explain how they relate to the plausibility of transitions between locations and cell states across time points.
@@
Spatial Smoothness. Tissues are well-organized systems. Within a microenvironment, neighboring cells respond to the same set of external mechanical stimuli and intercellular communication, which affect their states similarly and result in local smoothness in cell-type-specific expression. Due to tissue heterogeneity, we cannot assume a common reference coordinate frame across tissue samples or even slices at $t_{i}$ and $t_{j}$ at a larger scale. However, the same heterogeneity allows us to consider the spatial coherence and neighborhood consistency (Greenwald et al., 2024; Ceccarelli et al., 2025) as a proxy for relative cell localization, which cannot change significantly across short time intervals. Therefore the aggregate expression within the microenvironment of each cell can be used to quantify the transitional plausibility in consecutive time points.
@@
$\mathrm{SS}(c_{i},c_{j})$ $=\bigg\|\frac{1}{|\mathcal{N}_{r}(c_{i})|}\sum_{c\in\mathcal{N}_{r}(c_{i})}\bm{x}(c)$ (9)
$\qquad\qquad-\frac{1}{|\mathcal{N}_{r}(c_{j})|}\sum_{c\in\mathcal{N}_{r}(c_{j})}\bm{x}(c)\bigg\|_{2}^{2},$
where $\mathcal{N}_{r}(c_{i})=\{c:\|\bm{s}(c)-\bm{s}(c_{i})\|_{2}\leq r\}$ denotes the set of neighboring cells of $c_{i}$ in the same tissue slice, $|\mathcal{N}_{r}(c_{i})|$ stands for the cardinality of $\mathcal{N}_{r}(c_{i})$, and $\bm{x}(c)$ is the gene expression profile with respect to cell $c$.
@@
Specifically, we represent each cell $c_{i}$ by a feature vector $f_{\mathrm{LR}}\in\mathbb{R}^{p}$, where each entry corresponds to one of $p$ possible ligand–receptor pairs and encodes the extent of $c_{i}$’s participation in communication through that pair. The transitional plausibility $\mathrm{TP}(c_{i},c_{j})$ between cells in different tissue slices is higher when they exhibit similar ligand-receptor communication patterns $f_{\mathrm{LR}}$ (see Figure 8 for an illustration). Formally, we define $\mathrm{LR}(c_{i},c_{j})$, the dissimilarity between the ligand–receptor communication patterns in the microenvironments of cells $c_{i}$ and $c_{j}$, as:
$\mathrm{LR}(c_{i},c_{j})=\left\|\>f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{i}))-f_{\mathrm{LR}}(\mathcal{N}_{r}(c_{j}))\right\|_{2}^{2}.$ (10)
3.3 Flow Matching with Context-Aware OT Couplings
@@
Transitional Plausibility Matrix. To begin with, we create a sequence of transitional plausibility matrices (TPMs) to encode the biological priors for each pair of consecutive time points. Specifically, denote by $\mathbf{M}_{i,i+1}\in\mathbb{R}^{n_{i}\times n_{i+1}}$ the TPM with respect to the set of cells measured at time $t_{i}$ and at time $t_{i+1}$, with size $n_{i}$ and $n_{i+1}$ respectively, where the $(k,l)$-th entry of $\mathbf{M}_{i,i+1}$ indicates how plausibly the $k$-th cell measured at $t_{i}$ will evolve to the $l$-th cell measured at $t_{i+1}$, defined as follows:
$\big[\mathbf{M}_{i,i+1}\big]_{kl}=\lambda\cdot\mathrm{SS}$ $\left(c_{i}(k),c_{i+1}(l)\right)$ (11)
$+$ $(1-\lambda)\cdot\mathrm{LR}\left(c_{i}(k),c_{i+1}(l)\right).$
@@
Here, $\lambda\in[0,1]$ denotes a trade-off hyperparameter that balances the contribution of the spatial smoothness prior ($\mathrm{SS}$) and the ligand–receptor communication prior ($\mathrm{LR}$).
@@
Prior-Regularized OT Couplings. The transitional plausibility matrices capture our spatially informed prior over cell-cell transitions between consecutive time points, which can naturally be incorporated into the EOT formulation (Equation 7) to promote couplings that maintain the structural and functional properties of tissue organization. We propose two techniques for prior integration:
@@
Prior-Aware Cost Matrix (PACM). Consider the empirical counterpart of Equation 7 with respect to time $t_{i}$ and time $t_{i+1}$. Our first PACM approach incorporates the transitional plausibility matrix directly into the transport cost:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\;\langle\mathbf{\Pi},\mathbf{C}_{i,i+1}\rangle-\epsilon\sum_{k,l}\Pi_{kl}(\log\Pi_{kl}-1),$ (12)
@@
where the $(k,l)$-th entry of the new prior-aware cost matrix is $\mathbf{C}_{i,i+1}]_{kl}=\alpha\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}+(1-\alpha)[\mathbf{M}_{i,i+1}]_{kl}$, the transport plan $\Pi$ needs to satisfy the boundary conditions: $\sum_{l}\Pi_{kl}=1/n_{i}$ for any $k\in[n_{i}]$, and $\sum_{k}\Pi_{kl}=1/n_{i+1}$ for any $l\in[n_{i+1}]$, and $\alpha\in[0,1]$ controls the trade-off between the original Euclidean cost and the prior-aware cost derived from the transitional plausibility. If $\mathbf{M}_{i,i+1}]_{kl}$ is high, Equation 12 will impose a higher transport cost between the $k$-cell at time $i$ to the $j$-cell at time $i+1$. This observation aligns with our assumption that such transitions should be more biologically implausible.
@@
Prior-Aware Entropy Regularization (PAER). While the previous PACM approach penalizes couplings according to our spatial priors, it defines a different OT problem with a modified cost function. Consequently, the standard interpretation of OT as minimizing the transport energy between two transcriptomic distributions no longer holds. Since the scales of the pairwise distances often differ, normalization of the cost terms is required to enable meaningful comparison. This normalization, however, may result in couplings that deviate from their original counterparts (see Proposition 1 and Corollary 1 in Appendix D). Besides, selecting an appropriate $\alpha$ in Equation 12 introduces an additional layer of tuning, increasing computational overhead. Therefore, we propose the second PAER approach to integrate the biological priors without introducing additional hyperparameters:
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{i}\times n_{i+1}}}\sum_{k,l}\Pi_{kl}\|\bm{x}_{i}(k)-\bm{x}_{i+1}(l)\|_{2}^{2}$ (13)
@@
$\qquad\qquad-\epsilon\sum_{k,l}\underbrace{\Pi_{kl}\big(\log\big(\Pi_{kl}/[\widehat{\mathbf{M}}_{i,i+1}]_{kl}\big)-1\big)}_{\text{Prior-Aware Entropy Regularization}},$
@@
where $\widehat{\mathbf{M}}_{i,i+1}]_{kl}=\dfrac{\exp(-[\mathbf{M}_{i,i+1}]_{kl})}{\sum_{l}\exp(-[\mathbf{M}_{i,i+1}]_{kl})}$ denotes the $(k,l)$-th entry of the prior joint probability matrix induced by the originally constructed $\mathbf{M}_{i,i+1}$. Note that the lower the cost $\mathbf{M}_{i,i+1}]_{kl}$, the larger the entry $\widehat{\mathbf{M}}_{i,i+1}]_{kl}$, suggesting a higher plausibility of the transition from cell $k$ at $t_{i}$ to cell $l$ at $t_{i+1}$. The entropy regularization term in Equation 13 thus biases the learned transport plan toward the prior $\widehat{\mathbf{M}}_{i,i+1}$ rather than a uniform baseline, providing a soft mechanism for incorporating biological prior knowledge.
@@
ContextFlow. We apply the Sinkhorn algorithm (Cuturi, 2013) to solve the optimization problems defined in Equations 12 and 13 to obtain the spatial context-aware EOT couplings, and train the neural velocity vector field $u_{\theta}$ based on stochastic gradient descent by minimizing the multi-time generalization of Equation 3 with respect to conditionals $p_{t}(\bm{x}|\bm{z})$ and $u_{t}(\bm{x}|\bm{z})$ defined according to Equation 3.1.
The pseudocode for the proposed method, named Conditional Flow Matching with Context-Aware OT Couplings, is presented in Algorithm 1 in Appendix E.1. We also provide time complexity analysis in Appendix E.2. In particular, to apply the Sinkhorn algorithm to solve our PAER problem in Equation 13, we make use of the following theorem, a generalized result of Peyré et al. (2019).
Theorem 3.1.
$\min_{\mathbf{\Pi}\in\mathbb{R}^{n_{0}\times n_{1}}}\sum_{k,l}\Pi_{kl}C_{kl}-\epsilon\sum_{k,l}\Pi_{kl}\big(\log(\Pi_{kl}/{M}_{kl})-1\big),$
Theorem 3.1, proven in Appendix C, suggests a new Gibbs kernel $\mathbf{K}=\mathbf{M}\odot\exp(-\mathbf{C}/\epsilon)$, which combines both the transport cost and the prior joint probability matrices. When $\epsilon\rightarrow 0$, $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathbf{\Pi}_{\mathrm{ot}}^{*}$, thereby recovering the standard OT couplings in Equation 14. When $\epsilon\rightarrow\infty$, the optimal coupling $\mathbf{\Pi}^{*}_{\mathrm{CTF}}\rightarrow\mathrm{diag}(\bm{u})\cdot\mathbf{M}\cdot\mathrm{diag}(\bm{v})$, which corresponds to a plan that aligns with the prior defined by $\mathbf{M}$ rather than the independent couplings obtained with EOT defined by Equation 7. This has the same effect as constraining our transport plan through the proposed prior and, by extension, the flow. By varying the parameter $\epsilon$, we can thus efficiently optimize for a desirable coupling via the Sinkhorn algorithm.
4 Experiments
Datasets. We evaluate ContextFlow on three longitudinal spatial transcriptomics datasets: Axolotl Brain Regeneration (Wei et al., 2022), Mouse Embryo Organogenesis (Chen et al., 2022), and Liver Regeneration (Ben-Moshe et al., 2021). For all the evaluated datasets, the gene expression values are log-normalized, and we extract the top $50$ principal components (PCs) as feature vectors. The strength of ligand–receptor interactions in the microenvironment was inferred using spatially informed bivariate statistics implemented in LIANA+ (Dimitrov et al., 2024), where we applied the cosine similarity metric to gene expression profiles. Interaction evidence was aggregated using the consensus of multiple curated ligand–receptor resources, ensuring robustness of the inferred signals.
Baselines & Metrics. We benchmark ContextFlow using its two prior integration strategies—cost-regularized (CTF-C) and entropy-regularized (CTF-H)—against several existing flow matching baselines. As a non-spatial baseline, we include conditional flow matching (CFM), which uses only transcriptomic data with random couplings. We further compare against minibatch-OT flow matching (MOTFM), which leverages OT-derived couplings but does not incorporate spatial priors. For evaluation, we employed the 2-Wasserstein distance ($\mathcal{W}_{2}$), a commonly used OT-based metric, as well as MMD and Energy Distance for statistical fidelity. Furthermore, to assess the biological plausibility of our predicted dynamics, we evaluate them using a cell-type-weighted Wasserstein distance (Weighted $\mathcal{W}_{2}$), where the weights correspond to the relative frequency of each cell type in the dataset. Exact metric definitions are present in Appendix B.2. To account for randomness, all reported metrics are averaged across $10$ runs.